# TOAST: Fast and scalable auto-partitioning based on principled static analysis

### v1

@@ 1. Introduction
  • The other auto-sharding tool (TOAST). An automatic partitioning tool that integrates our analysis with a Monte-Carlo Tree Search agent (Browne et al., 2012). The analysis guides the agent by (1) augmenting the search space with conflict resolution actions and (2) pruning the space by identifying dimensions that must be sharded together. This design enables the efficient discovery of complex resharding strategies (Section 4).
@@ 2.1. Partitioning of ML models
In the context of ML models, it is a common abstraction to organize the available accelerator devices (e.g. GPUs, TPUs) into a (logical) mesh, i.e. into an $n$-dimensional lattice that is spanned by $n$ device axes (cf. (Lepikhin et al., 2020; Xu et al., 2021; Shazeer et al., 2018; Alabed et al., 2024)). In this section, we assume that devices are organized in a 2-dim. mesh, spanned by axes b and m. If there are $b$ devices along axis b and $m$ devices along m, then the whole mesh consists of $b\cdot m$ devices.
Partitioning an ML model is the process of preparing it for execution on a mesh of devices. This process involves distributing the model’s tensors across the devices as smaller chunks, shards. A shard is created by slicing a tensor along one of its dimensions and assigning it to devices along a corresponding axis of the device mesh, as illustrated in Figure 1.
@@
To illustrate these concepts, we look at the simplified MLP (multi-layer perceptron) in Figure 2. There are two matrix multiplications, and a non-linear operation (ReLU) in between. The inputs, or samples, are passed in as x, and the model parameters are given as arguments w1, w2.
We observe that all operations in the body of mlp act as maps across the tensor dimensions highlighted in yellow: 256. Thus, all operations can be executed in parallel on shards of tensors along these dimensions. Figure 2 uses the annotation {b} to indicate that these dimensions have been sharded along axis b. Although the body of mlp appears unchanged, its operations now execute on smaller tensors, specifically the shards corresponding to the full tensors shown in Figure 2. Figure 2 is the batch-partitioned version of the mlp function. The function now operates on a shard of the input samples x, known as a batch. This listing shows the device-local code for mlp, where each device initially holds one such shard of x. During execution, each device computes shards of intermediate tensors. Consequently, the returned tensor w on each device represents a shard of the full output tensor from Figure 2. Our batch partitioning of mlp requires no communication between devices.
@@ 2.2. Partitioning, algorithmically
We have demonstrated batch and Megatron partitioning strategies on our mlp example. Each strategy requires sharding a specific set of tensors along particular dimensions, corresponding to the colored dimensions in Figure 2. In this manual approach, the correct dimensions were identified by inspecting the operations in the mlp with the goal of minimizing communication.

### v2

@@ 1. Introduction
  • The other auto-sharding tool (TOAST). An automatic partitioning tool that integrates our analysis with a Monte-Carlo Tree Search (MCTS) agent (Browne et al., 2012). The analysis guides the agent by (1) augmenting the search space with conflict resolution actions and (2) pruning the space by identifying dimensions that must be sharded together. This design enables the efficient discovery of complex resharding strategies (Section 4).
@@ 2.1. Partitioning of ML models
When scaling ML models, it is a common abstraction to organize available accelerator devices (e.g., GPUs, TPUs) into a logical mesh. This mesh is an $n$-dimensional lattice spanned by $n$ device axes (Lepikhin et al., 2020; Xu et al., 2021; Shazeer et al., 2018; Alabed et al., 2024). In this section, we assume that devices are organized in a 2-dim. mesh, spanned by axes b and m. If there are $b$ devices along axis b and $m$ devices along m, then the whole mesh consists of $b\cdot m$ devices.
Partitioning a ML model is the process of preparing it for execution on a mesh of devices. It involves distributing the model’s tensors across the devices as smaller chunks, known as shards. A shard is created by slicing a tensor along one of its dimensions and assigning it to devices along a corresponding axis of the device mesh, as illustrated in Figure 1.
@@
The simplified multi-layer perceptron (mlp) in Figure 2 illustrates these concepts. The model consists of two matrix multiplications separated by a non-linear operation (ReLU). It takes the input samples, x, and model parameters, w1 and w2, as arguments. All operations in the body of mlp are mapped across the tensor dimension highlighted in yellow: 256. Consequently, these operations can be executed in parallel on tensor shards created along this dimension. Figure 2 uses the annotation {b} to indicate that these dimensions have been sharded along axis b. Although the body of mlp appears unchanged, its operations now execute on smaller tensors, specifically the shards corresponding to the full tensors shown in Figure 2. Figure 2 is the batch-partitioned version of the mlp function. The function now operates on a shard of the input samples x, known as a batch. This listing shows the device-local code for mlp, where each device initially holds one such shard of x. During execution, each device computes shards of intermediate tensors. Consequently, the returned tensor w on each device represents a shard of the full output tensor from Figure 2. Our batch partitioning of mlp requires no communication between devices.
@@ 2.2. Partitioning, algorithmically
We demonstrated batch and Megatron partitioning strategies on our mlp example. Each strategy requires sharding a specific set of tensors along particular dimensions, corresponding to the colored dimensions in Figure 2. In manual approach, the dimensions were identified by inspecting the operations in the mlp to minimize communication.